Lagard's stock had a required return of 11.75% last year, when the risk-free rate was 5.50% and the market risk premium was 4.75%. Then an increase in investor risk aversion caused the market risk premium to rise by 1.5%. The risk-free rate and the firm's beta remain unchanged. What is the company's new required rate of return? (Hint: First calculate the beta, then find the required return.)
Answer #1

required return=risk free rate+Beta*market risk premium

Currently:

11.75=5.5+Beta*4.75

Beta=(11.75-5.5)/4.75

=1.31578947(Approx)

New market risk premium=4.75+1.5=6.25%

Hence new required rate=5.5+(1.31578947*6.25)

=13.72%(Approx).